OpenStudy (anonymous):

The length of a rectangle is 8 inches and the width is x inches. What is the perimeter, in inches? A. 16 + 2x B. 10 + 2x C. 16 + x D. 10 + x

OpenStudy (acxbox22):

perimeter=2(l+w) p=2(x+8) p=2x+16